OpenStudy (anonymous):

my question is f(X + 1) = . the rule is f(X) =3X+2. i got - 3(X + 1) +2 then i got - 3X +5 can someone explain what to do from here, im stuck

OpenStudy (anonymous):

thats it, your done, unless they are asking you to solve for x

hero (hero):

3(x+ 1) + 2 = 3x + 5

OpenStudy (anonymous):

when you sub in an expression, you often get an expression as the simplified form

OpenStudy (anonymous):

thx very much i feel real dumb right about now but its watever , i guess <3 ; ) <3

OpenStudy (anonymous):

they r asking mi to solve for x

OpenStudy (anonymous):

you've got f(x+1)=3x+5 but this is not an equation. you can solve for x if you let y=0, where y=f(x+1). is that what they mean? solve 3x+5=0

OpenStudy (anonymous):

3x+5=0 3x=-5 x=-5/3 but i am just guess that that i what they mean